OpenStudy (anonymous):

An environmental agency studying the effects of drought on a lake found that the water level decreased at a constant rate during the first few months of the drought. When the agency began its research, the depth of the lake was 346 feet. The table represents the agency’s susequent observations about the lake's depth. Which equation represents the depth of the lake, y, based on the number of weeks passed, x? y = -0.6x + 346 y = -1.2x + 346 y = -0.6x y = -1.2x

OpenStudy (anonymous):

? An environmental agency studying the effects of drought on a lake found that the water level decreased at a constant rate during the first few months of the drought. When the agency began its research, the depth of the lake was 346 feet. The table represents the agency’s susequent observations about the lake's depth. Which equation represents the depth of the lake, y, based on the number of weeks passed, x? y = -0.6x + 346 y = -1.2x + 346 y = -0.6x y = -1.2x Done Time (weeks) Lake Depth (feet) 0 346.0 2 344.8 5 343.0 7 341.8 10 340.0 12 338.8

OpenStudy (anonymous):

hello again

OpenStudy (anonymous):

hello

OpenStudy (camerondoherty):

Answer: a)y = -0.6x + 346

OpenStudy (anonymous):

i just tryed it i'm not that good in math and i got a too

OpenStudy (camerondoherty):

did you get it @Gallardo_kidd ???

OpenStudy (anonymous):

hold on

OpenStudy (camerondoherty):

(-_-)

OpenStudy (anonymous):

it was right thank you do you have a facebook

OpenStudy (anonymous):

During a construction project, engineers used explosives to excavate 140 feet of tunnel into a mountain. But because of time constraints and environmental concerns, they brought in a tunnel boring machine (TBM) to excavate the rest of the tunnel. The data table lists some observations an engineer made about the length of the tunnel after the TBM was introduced. Which equation represents the length of the completed tunnel based on the number of days since TBM was introduced? y = 45x + 140 y = 405x + 140 y = 68.33x + 140 y = 25x + 140

OpenStudy (anonymous):

? During a construction project, engineers used explosives to excavate 140 feet of tunnel into a mountain. But because of time constraints and environmental concerns, they brought in a tunnel boring machine (TBM) to excavate the rest of the tunnel. The data table lists some observations an engineer made about the length of the tunnel after the TBM was introduced. Which equation represents the length of the completed tunnel based on the number of days since TBM was introduced? y = 45x + 140 y = 405x + 140 y = 68.33x + 140 y = 25x + 140 Done TBM Use (days) Tunnel Length (feet) 6 410 15 815 20 1,040 28 1,400 33 1,625 38 1,850 45 2,165

OpenStudy (camerondoherty):

can i have a medal plz? :)

OpenStudy (camerondoherty):

Thanks and hold on lemme see

OpenStudy (anonymous):

your welcome

OpenStudy (camerondoherty):

ask a new question instead of putting 2 questions in 1 thing

OpenStudy (anonymous):

thats was one question